please tell me how to solve
what is the difference between the sum of the prime factors of 18 and the sum of the prime factor of 25?

Answer by praseenakos@yahoo.com(507)

QUESTION

what is the difference between the sum of the prime factors of 18 and the sum of the prime factor of 25?

ANSWER:


Prime Numbers.


A prime number is a positive integer that has exactly two positive integer factors, 1 and itself.



For example, if we list the factors of 28, we have 1, 2, 4, 7, 14, and 28.



That's six factors. If we list the factors of 29, we only have 1 and 29.



That's two factors. So we say that 29 is a prime number, but 28 isn't.



Note that the definition of a prime number doesn't allow 1 to be a prime number:


1 only has one factor, namely 1.


Prime numbers have exactly two factors, not "at most two" or anything like that. When a number has more than two factors it is called a composite number.



Here are the first few prime numbers:



2, 3, 5, 7, 11, 13, 17, 19, 23, 29, 31, 37, 41, 43, 47, 53, 59, 61, 67, 71, 73, 79, 83, 89, 97, 101, 103, 107, 109, 113, 127, 131, 137, 139, 149, 151, 157, 163, 167, 173, 179, 181, 191, 193, 197, 199, etc.


Now let's come back to your question.


List out prime factors of 18 and 25


Factors of 18 = 1, 2, 3, 6, 9, and 18


Among these prime numbers are 2 and 3, all others are composite.


Factors of 25 = 1, 5, 25.



Among these only 5 is prime.


Now sum of prime factors of 18 = 2+3 = 5


sum of prime factors of 25 = 5


Their difference is = 5-5=0


So your answer is 0



Hope you understood.

Regards.
praseenakos@yahoo.co.in